# Corruption-Robust Linear Bandits: Minimax Optimality and Gap-Dependent Misspecification

### v1

@@ 5.2 $C$ bound in Adversarial Linear Bandits
For SCRiBLe and CEW, the work by Lee et al., (2020) and Zimmert and Lattimore, (2022) developed techniques that incorporate bonus terms to get high probability regret bounds. The bonus terms introduced by Zimmert and Lattimore, (2022) is similar to that discussed in Eq. (7), which only allows us to get a $C_{\textsf{sq}}$ bound. The bonus terms introduced by Lee et al., (2020) allows us to obtain a $C$ bound, but the overhead introduced by the bonus terms is much larger, resulting in a highly sub-optimal regret bound. Indeed, as shown in Appendix H, adopting their bonus construction results in an additional regret of $d^{\frac{5}{2}}C$. With several attempts, we are only able to obtain the tight corruption dependency $dC$ using the bonus in Section 5.1. To use that bonus, however, it is necessary to lift the problem to $(d+1)^{2}$-dimensional space. Unfortunately, existing SCRiBLe and CEW algorithms only operate in the original $d$-dimensional space, and as discussed above, we need them to ensure $\|a_{t}\|_{\Sigma_{t}^{-1}}\leq\operatorname{poly}(d)$.
@@
Note that CEW requires the assumption that the feasible set is a convex body with non-zero volume, but the effective dimension of ${\mathcal{H}}$ is strictly smaller than $(d+1)^{2}$ and thus have zero volume in $\mathbb{R}^{(d+1)^{2}}$. To correctly write the algorithm, we introduce an invertible linear transformation $\phi:\mathbb{R}^{(d+1)^{2}}\to\mathbb{R}^{m}$ that maps an $(d+1)^{2}$-dimensional action set ${\mathcal{H}}$ to an $m$-dimensional one, where $m$ is the effective dimension of ${\mathcal{H}}$. In Appendix LABEL:app:_continuous_ew_for_low_dim, we formally define this $\phi$. The algorithm uses $\phi$ to map all lifted actions and reward estimators from $\mathbb{R}^{(d+1)\times(d+1)}$ to $\mathbb{R}^{m}$.
The exponential weights runs over the space of $\phi({\mathcal{H}})$ (see Eq. (10)). A point $h\in\phi({\mathcal{H}})$ sampled from the exponential weights can be linearly mapped to an action $a\in{\mathcal{A}}$ according to Eq. (11). We use $q_{t}^{\prime}$ to denote the exponential weight distribution in $\phi({\mathcal{H}})$, and use $p_{t}$ to denote the corresponding distribution in ${\mathcal{A}}$. Instead of sampling $a_{t}$ from $p_{t}$, we sample it through rejection sampling that rejects samples with $\|a_{t}\|_{\Sigma_{t}^{-1}}>\tilde{\Theta}(\sqrt{d})$ (Eq. (12)). This technique was developed by Ito et al., (2020), and this guarantees $\|a_{t}\|_{\Sigma_{t}^{-1}}\leq\tilde{\mathcal{O}}(\sqrt{d})$—which is our goal as discussed in Eq. (9)—while keeping the clipped distribution $\tilde{p}_{t}$ close enough to the original distribution $p_{t}$. This last property heavily relies on the log-concavity of the exponential weight distribution (Ito et al.,, 2020). The definition of the bonus term is similar to that in Algorithm 2 (Eq. (13)). The construction of the reward estimator (Line 3) and the way of lifting (Eq. (10)) are also similar to those in Algorithm 2. Again, we adopt the fixed-point formulation where the calculation of the policy at time $t$ involves the bonus at time $t$, which, in turn, depends on the policy at time $t$. It is unlikely that this algorithm can be polynomial time. As a remedy, we provide a polynomial time algorithm (Algorithm 5) in Appendix H with a much worse regret bound of $\tilde{\mathcal{O}}(d^{3}\sqrt{T}+d^{\nicefrac{{5}}{{2}}}C)$. The regret guarantee of Algorithm 3 is given in the following theorem.
@@ 6.1 Optimal Misspecification Rate for Linear Bandits
Theorem 6.2 relies on a specialized algorithm for the gap-dependent misspecification setting, and improves on the best-known bound for gap-dependent misspecification in linear bandits, which requires $\rho\leq\widetilde{\mathcal{O}}(\frac{1}{d})$ (Liu et al., 2023a, ). Moreover, for $\rho>c_{T}\frac{1}{\sqrt{d}}$ for some logarithmic term $c_{T}$, adapting the lower-bound instance from Lattimore et al., (2020), we show that achieving sub-linear regret is not possible (Theorem I.2). These results jointly show that $\rho\approx\frac{1}{\sqrt{d}}$ is the best $\rho$ we can hope for. This disproves the conjecture of Liu et al., 2023a that $\rho=\Theta(1)$ is possible.
@@ 6.2 Gap-Dependent Misspecification in Reinforcement Learning
Theorem 6.1 is a corollary of a more general result, Theorem J.1, which applies to misspecified reinforcement learning, where we there assume a generalized notion of gap-dependent misspecification: for each policy $\pi$, $\mathbb{E}^{\mathcal{M}^{\star},\pi}[\sum_{h=1}^{H}\epsilon^{\mathrm{mis}}_{h}(s_{h},a_{h})]\leq\rho\cdot(V_{0}^{\star}-V_{0}^{\pi})$, for $V_{0}^{\pi}$ the expected reward of policy $\pi$, and $\epsilon^{\mathrm{mis}}_{h}(s,a)$ a measure of the misspecification at step $h$, state $s$, and action $a$. To illustrate this general reduction, we consider the following setting, a generalization of linear MDPs (Jin et al.,, 2020).
@@ Appendix D Proof of Proposition 1
By the strengthened Freedman’s inequality (Lemma K.3), with probability at least $1-\delta$,
@@
$\left\langle\boldsymbol{H}-\boldsymbol{H}_{t},\boldsymbol{\widehat{\gamma}}_{t}\right\rangle-\frac{D_{G}(\boldsymbol{H},\boldsymbol{H}_{t})}{2\eta}$ $\leq\left\langle\boldsymbol{H}-\boldsymbol{H}_{t},\boldsymbol{\widehat{\gamma}}_{t}\right\rangle-\frac{\|x_{t}-h\|_{\text{\rm Cov}(p_{t})^{-1}}^{2}}{2\eta}$ (Lemma K.1)
@@ 6       Compute
where the last inequality is due to $\Sigma_{t}^{-1}\succeq B_{t}-B_{t-1}$. Since $\eta\leq\frac{1}{16\sqrt{d}\alpha}$, by Lemma K.2, with probability of at least $1-\delta$, we have
@@
Appendix H Computationally Efficient Algorithm for Adversarial $C$ Bound

### v2

@@ 5.2 $C$ bound in Adversarial Linear Bandits
For SCRiBLe and CEW, the work by Lee et al., (2020) and Zimmert and Lattimore, (2022) developed techniques that incorporate bonus terms to get high probability regret bounds. The bonus terms introduced by Zimmert and Lattimore, (2022) is similar to that discussed in Eq. (7), which only allows us to get a $C_{\textsf{sq}}$ bound. The bonus terms introduced by Lee et al., (2020) allows us to obtain a $C$ bound, but the overhead introduced by the bonus terms is much larger, resulting in a highly sub-optimal regret bound. Indeed, as shown in Appendix J, adopting their bonus construction results in an additional regret of $d^{\frac{5}{2}}C$. With several attempts, we are only able to obtain the tight corruption dependency $dC$ using the bonus in Section 5.1. To use that bonus, however, it is necessary to lift the problem to $(d+1)^{2}$-dimensional space. Unfortunately, existing SCRiBLe and CEW algorithms only operate in the original $d$-dimensional space, and as discussed above, we need them to ensure $\|a_{t}\|_{\Sigma_{t}^{-1}}\leq\operatorname{poly}(d)$.
@@
Note that CEW requires the assumption that the feasible set is a convex body with non-zero volume, but the effective dimension of ${\mathcal{H}}$ is strictly smaller than $(d+1)^{2}$ and thus have zero volume in $\mathbb{R}^{(d+1)^{2}}$. To correctly write the algorithm, we introduce an invertible linear transformation $\phi:\mathbb{R}^{(d+1)^{2}}\to\mathbb{R}^{m}$ that maps an $(d+1)^{2}$-dimensional action set ${\mathcal{H}}$ to an $m$-dimensional one, where $m$ is the effective dimension of ${\mathcal{H}}$. In Appendix I, we formally define this $\phi$. The algorithm uses $\phi$ to map all lifted actions and reward estimators from $\mathbb{R}^{(d+1)\times(d+1)}$ to $\mathbb{R}^{m}$.
The exponential weights runs over the space of $\phi({\mathcal{H}})$ (see Eq. (10)). A point $h\in\phi({\mathcal{H}})$ sampled from the exponential weights can be linearly mapped to an action $a\in{\mathcal{A}}$ according to Eq. (11). We use $q_{t}^{\prime}$ to denote the exponential weight distribution in $\phi({\mathcal{H}})$, and use $p_{t}$ to denote the corresponding distribution in ${\mathcal{A}}$. Instead of sampling $a_{t}$ from $p_{t}$, we sample it through rejection sampling that rejects samples with $\|a_{t}\|_{\Sigma_{t}^{-1}}>\tilde{\Theta}(\sqrt{d})$ (Eq. (12)). This technique was developed by Ito et al., (2020), and this guarantees $\|a_{t}\|_{\Sigma_{t}^{-1}}\leq\tilde{\mathcal{O}}(\sqrt{d})$—which is our goal as discussed in Eq. (9)—while keeping the clipped distribution $\tilde{p}_{t}$ close enough to the original distribution $p_{t}$. This last property heavily relies on the log-concavity of the exponential weight distribution (Ito et al.,, 2020). The definition of the bonus term is similar to that in Algorithm 2 (Eq. (13)). The construction of the reward estimator (Line 3) and the way of lifting (Eq. (10)) are also similar to those in Algorithm 2. Again, we adopt the fixed-point formulation where the calculation of the policy at time $t$ involves the bonus at time $t$, which, in turn, depends on the policy at time $t$. It is unlikely that this algorithm can be polynomial time. As a remedy, we provide a polynomial time algorithm (Algorithm 6) in Appendix J with a much worse regret bound of $\tilde{\mathcal{O}}(d^{3}\sqrt{T}+d^{\nicefrac{{5}}{{2}}}C)$. The regret guarantee of Algorithm 3 is given in the following theorem.
@@ 6.1 Optimal Misspecification Rate for Linear Bandits
Theorem 6.2 relies on a specialized algorithm for the gap-dependent misspecification setting, and improves on the best-known bound for gap-dependent misspecification in linear bandits, which requires $\rho\leq\widetilde{\mathcal{O}}(\frac{1}{d})$ (Liu et al., 2023a, ). Moreover, for $\rho>c_{T}\frac{1}{\sqrt{d}}$ for some logarithmic term $c_{T}$, adapting the lower-bound instance from Lattimore et al., (2020), we show that achieving sub-linear regret is not possible (Theorem K.2). These results jointly show that $\rho\approx\frac{1}{\sqrt{d}}$ is the best $\rho$ we can hope for. This disproves the conjecture of Liu et al., 2023a that $\rho=\Theta(1)$ is possible.
@@ 6.2 Gap-Dependent Misspecification in Reinforcement Learning
Theorem 6.1 is a corollary of a more general result, Theorem L.1, which applies to misspecified reinforcement learning, where we there assume a generalized notion of gap-dependent misspecification: for each policy $\pi$, $\mathbb{E}^{\mathcal{M}^{\star},\pi}[\sum_{h=1}^{H}\epsilon^{\mathrm{mis}}_{h}(s_{h},a_{h})]\leq\rho\cdot(V_{0}^{\star}-V_{0}^{\pi})$, for $V_{0}^{\pi}$ the expected reward of policy $\pi$, and $\epsilon^{\mathrm{mis}}_{h}(s,a)$ a measure of the misspecification at step $h$, state $s$, and action $a$. To illustrate this general reduction, we consider the following setting, a generalization of linear MDPs (Jin et al.,, 2020).
@@ Appendix D Proof of Proposition 1
By the strengthened Freedman’s inequality (Lemma M.3), with probability at least $1-\delta$,
@@
$\left\langle\boldsymbol{H}-\boldsymbol{H}_{t},\boldsymbol{\widehat{\gamma}}_{t}\right\rangle-\frac{D_{G}(\boldsymbol{H},\boldsymbol{H}_{t})}{2\eta}$ $\leq\left\langle\boldsymbol{H}-\boldsymbol{H}_{t},\boldsymbol{\widehat{\gamma}}_{t}\right\rangle-\frac{\|x_{t}-h\|_{\text{\rm Cov}(p_{t})^{-1}}^{2}}{2\eta}$ (Lemma M.1)
@@ 6       Compute
where the last inequality is due to $\Sigma_{t}^{-1}\succeq B_{t}-B_{t-1}$. Since $\eta\leq\frac{1}{16\sqrt{d}\alpha}$, by Lemma M.2, with probability of at least $1-\delta$, we have
@@
Appendix H Proof of Theorem 5.2
Similar to before, we define
$\boldsymbol{\widehat{\gamma}}_{t}$ $=\begin{bmatrix}0&\frac{1}{2}\widehat{\theta}_{t}\\ \frac{1}{2}\widehat{\theta}_{t}^{\top}&0\end{bmatrix},\qquad\boldsymbol{D}_{t}=\begin{bmatrix}\alpha B_{t}-\alpha B_{t-1}&0\\ 0&0\end{bmatrix},$
and $x_{t}=\mathbb{E}_{a\sim p_{t}}[a]$, $\tilde{x}_{t}=\mathbb{E}_{a\sim\tilde{p}_{t}}[a]$. We perform the regret decomposition as the following.
$\textup{Reg}_{T}=\sum_{t=1}^{T}\left\langle{u-a_{t},\theta_{t}}\right\rangle$
$=\sum_{t=1}^{T}\left\langle{u-x_{t},\theta_{t}}\right\rangle+\sum_{t=1}^{T}\left\langle{x_{t}-\tilde{x}_{t},\theta_{t}}\right\rangle+\sum_{t=1}^{T}\left\langle{\tilde{x}_{t}-a_{t},\theta_{t}}\right\rangle$
$=\sum_{t=1}^{T}\left\langle{u-x_{t},\theta_{t}}\right\rangle+\mathcal{O}\big{(}\gamma T+\sqrt{T\log(1/\delta)}\big{)}$
$=\underbrace{\sum_{t=1}^{T}\left\langle{u-x_{t},\theta_{t}-\mathbb{E}_{t}[\widehat{\theta}_{t}]}\right\rangle}_{\textbf{{Bias}}}+\underbrace{\sum_{t=1}^{T}\left\langle{u-x_{t},\mathbb{E}_{t}[\widehat{\theta}_{t}]-\widehat{\theta}_{t}}\right\rangle}_{\textbf{{Deviation}}}+\underbrace{\sum_{t=1}^{T}\left\langle{u-x_{t},\widehat{\theta}_{t}}\right\rangle}_{\textbf{{FTRL}}}+\mathcal{O}\big{(}\gamma T+\sqrt{T\log(1/\delta)}\big{)}.$ (23)
The FTRL term can be further bounded as the following.
FTRL
$=\sum_{t=1}^{T}\mathbb{E}_{a\sim p_{t}}\left[\left\langle{u-a,\widehat{\theta}_{t}}\right\rangle\right]$
$=\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{H}\sim q_{t}}\left[\left\langle{\boldsymbol{U}-\boldsymbol{H},\boldsymbol{\widehat{\gamma}}_{t}}\right\rangle\right]$
$\leq\frac{d^{2}\log T}{\eta}+\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{H}\sim q_{t}}\left[\exp\left(\eta\left\langle{\boldsymbol{H},\boldsymbol{\widehat{\gamma}}_{t}}\right\rangle\right)-\eta\left\langle{\boldsymbol{H},\boldsymbol{\widehat{\gamma}}_{t}}\right\rangle-1\right]+\sum_{t=1}^{T}\mathbb{E}_{\boldsymbol{H}\sim q_{t}}\left[\left\langle{\boldsymbol{H}-\boldsymbol{U},\boldsymbol{D}_{t}}\right\rangle\right]$ (by Theorem I.2)
$=\frac{d^{2}\log T}{\eta}+\underbrace{\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{a\sim p_{t}}\left[\exp\left(\eta\left\langle{a,\widehat{\theta}_{t}}\right\rangle\right)-\eta\left\langle{a,\widehat{\theta}_{t}}\right\rangle-1\right]}_{\textbf{{Stability}}}+\underbrace{\alpha\sum_{t=1}^{T}\mathbb{E}_{a\sim p_{t}}\left[\|a\|_{B_{t}-B_{t-1}}^{2}\right]-\alpha\sum_{t=1}^{T}\|u\|_{B_{T}}^{2}}_{\textbf{{Bonus}}}.$ (24)
[…]
Lemma H.1.
Proof.
Bias $=\sum_{t=1}^{T}\left\langle u-x_{t},\theta_{t}-\mathbb{E}_{t}[\widehat{\theta}_{t}]\right\rangle$
$=\sum_{t=1}^{T}\left\langle u-x_{t},\theta_{t}-\widetilde{\Sigma}_{t}^{-1}\mathbb{E}_{t}\left[a_{t}a_{t}^{\top}\right]\theta_{t}+\widetilde{\Sigma}_{t}^{-1}\mathbb{E}_{t}\left[a_{t}\epsilon_{t}(a_{t})\right]\right\rangle$
$=\gamma\sum_{t=1}^{T}\left\langle u-x_{t},\widetilde{\Sigma}_{t}^{-1}\theta_{t}\right\rangle+\sum_{t=1}^{T}\left\langle u-x_{t},-\widetilde{\Sigma}_{t}^{-1}\mathbb{E}_{t}\left[a_{t}\epsilon_{t}(a_{t})\right]\right\rangle$ ($\widetilde{\Sigma}_{t}=\gamma I+\mathbb{E}_{t}[a_{t}a_{t}^{\top}]$)
$\leq\gamma\sum_{t=1}^{T}\left\|x_{t}-u\right\|_{\widetilde{\Sigma}_{t}^{-1}}\|\theta_{t}\|_{\widetilde{\Sigma}_{t}^{-1}}+\sum_{t=1}^{T}\left\|x_{t}-u\right\|_{\widetilde{\Sigma}_{t}^{-1}}\mathbb{E}_{t}\left[\left\|a_{t}\right\|_{\widetilde{\Sigma}_{t}^{-1}}|\epsilon_{t}(a_{t})|\right]$
$\leq\left(\max_{t}\left\|x_{t}-u\right\|_{\widetilde{\Sigma}_{t}^{-1}}\right)\left(\gamma\sqrt{\frac{d}{\gamma}}T+\sum_{t=1}^{T}\mathbb{E}_{t}\left[\left\|a_{t}\right\|_{\widetilde{\Sigma}_{t}^{-1}}|\epsilon_{t}(a_{t})|\right]\right)$ ($\widetilde{\Sigma}_{t}\succeq\gamma I$ and $\|\theta_{t}\|_{2}\leq\sqrt{d}$)
$\leq\left(\max_{t}\left\|x_{t}-u\right\|_{\widetilde{\Sigma}_{t}^{-1}}\right)\left(\sqrt{d\gamma}T+\sqrt{d}\beta\sum_{t=1}^{T}\mathbb{E}_{t}\left[|\epsilon_{t}(a_{t})|\right]\right)$
$\leq\left(\max_{t}\left\|x_{t}-u\right\|_{\widetilde{\Sigma}_{t}^{-1}}\right)\left(\sqrt{d\gamma}T+2\sqrt{T\log(1/\delta)}+\sqrt{d}\beta\sum_{t=1}^{T}|\epsilon_{t}(a_{t})|\right).$ (Azuma’s inequality)
Lemma H.2.
Proof.
Notice that
$\left|\langle u-x_{t},\widehat{\theta}_{t}\rangle\right|$ $\leq\left|(u-x_{t})^{\top}\widetilde{\Sigma}_{t}^{-1}a_{t}\right|$
$\leq\|u-x_{t}\|_{\widetilde{\Sigma}_{t}^{-1}}\|a_{t}\|_{\widetilde{\Sigma}_{t}^{-1}}$
$\leq\sqrt{d}\beta\|u-x_{t}\|_{\widetilde{\Sigma}_{t}^{-1}}.$
By the strengthened Freedman’s inequality (Lemma M.3), with probability at least $1-\delta$,
$\leq\mathcal{O}\left(3\sqrt{\sum_{t=1}^{T}\mathbb{E}_{t}\left[\left\langle{u-x_{t},\widehat{\theta}_{t}}\right\rangle^{2}\right]\log(T^{d}/\delta)}+2\sqrt{d}\beta\max_{t}\|u-x_{t}\|_{\widetilde{\Sigma}_{t}^{-1}}\log(T^{d}/\delta)\right)$
$\leq\mathcal{O}\left(\max_{t}\|u-x_{t}\|_{\widetilde{\Sigma}_{t}^{-1}}d\beta\sqrt{T}\log(T/\delta)\right).$ (using the assumption $d\leq T$)
Lemma H.3.
$\textbf{{Bonus}}\leq 3\alpha d\log(T)-\alpha\max_{t}\|u\|_{\widetilde{\Sigma}_{t}^{-1}}^{2}.$
Proof.
The proof the same as in the logdet case. See the proof of Lemma F.6.
Lemma H.4.
$\textbf{{Stability}}\leq\mathcal{O}(\eta dT\log^{2}T).$
Proof.
$\textbf{{Stability}}=\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{a\sim p_{t}}\left[\exp\left(\eta\left\langle{a,\widehat{\theta}_{t}}\right\rangle\right)-\eta\left\langle{a,\widehat{\theta}_{t}}\right\rangle-1\right]$
Since $q_{t}^{\prime}$ is a log-concave distribution, so are $q_{t}$ and $p_{t}$, which further implies that $\eta\left\langle{a,\widehat{\theta}_{t}}\right\rangle$ follows a log-concave distribution. Furthermore,
$\mathbb{E}_{a\sim p_{t}}\left[\eta^{2}\left\langle{a,\widehat{\theta}_{t}}\right\rangle^{2}\right]\leq\mathbb{E}_{a\sim p_{t}}\left[\eta^{2}a_{t}^{\top}\widetilde{\Sigma}_{t}^{-1}aa^{\top}\widetilde{\Sigma}_{t}^{-1}a_{t}\right]\leq 2\eta^{2}\|a_{t}\|_{\widetilde{\Sigma}_{t}^{-1}}^{2}\leq 2\eta^{2}d\beta^{2}\leq\frac{1}{100},$
where we use Lemma J.2 in the second-last inequality . By Lemma 6 of Ito et al., (2020), we have
$\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{a\sim p_{t}}\left[\exp\left(\eta\left\langle{a,\widehat{\theta}_{t}}\right\rangle\right)-\eta\left\langle{a,\widehat{\theta}_{t}}\right\rangle-1\right]$ $\leq\eta\sum_{t=1}^{T}\mathbb{E}_{a\sim p_{t}}\left[\left\langle{a,\widehat{\theta}_{t}}\right\rangle^{2}\right]\leq 2\eta\sum_{t=1}^{T}\|a_{t}\|^{2}_{\widetilde{\Sigma}_{t}^{-1}}\leq 2\eta\beta^{2}dT.$
Proof of Theorem 5.2.
Combining Eq. (23), Eq. (24), and Lemma H.1, Lemma H.2, Lemma H.3, Lemma H.4, we see that the regret is bounded by
$\tilde{\mathcal{O}}\left(\frac{d^{2}}{\eta}+\eta dT+\max_{t}\|u-x_{t}\|_{\widetilde{\Sigma}_{t}^{-1}}(d\sqrt{T}+\sqrt{d}C)+dC+\alpha d\right)-\alpha\|u\|^{2}_{B_{T}}$
$\leq\tilde{\mathcal{O}}\left(\frac{d^{2}}{\eta}+\eta dT+\alpha d\right)+\frac{d^{2}T+dC^{2}}{\alpha}$ (AM-GM inequality)
Choosing optimal $\alpha$ and $\eta$ leads to $\tilde{\mathcal{O}}(\sqrt{d^{3}T}+dC)$. ∎
Appendix I Dimension Reduction for Continuous Exponential Weights
First, the intrinsic dimension of ${\mathcal{X}}$ can be defined as the following:
Definition 1.
The intrinsic dimension of ${\mathcal{X}}$ is defined as
$\dim({\mathcal{X}})=\dim\left(\operatorname{span}\left({\mathcal{X}}-{\mathcal{X}}\right)\right),$
where ${\mathcal{X}}-{\mathcal{X}}\triangleq\left\{x-x^{\prime}:x,x^{\prime}\in{\mathcal{X}}\right\}$.
A convex region ${\mathcal{X}}\subset\mathbb{R}^{n}$ can be translated and rotated so that it entirely lies in $\mathbb{R}^{m}$ where $m=\dim({\mathcal{X}})$ and has non-zero volume in $\mathbb{R}^{m}$. We more precisely define this transformation below.
Definition 2.
Let ${\mathcal{X}}\subset\mathbb{R}^{n}$ be a convex region with $\dim({\mathcal{X}})=m$. We define $\phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ as the following linear transformation:
$\phi(x)\triangleq ZMx,$
where $M\in\mathbb{R}^{n\times n}$ is a rotation matrix (i.e., orthogonal matrix) such that for any $v\in{\mathcal{X}}-{\mathcal{X}}$, $Mv$ has non-zero elements only in the first $m$ coordinates (this is always possible by the definition of $\dim({\mathcal{X}})$ in Definition 1), and
$Z=\begin{bmatrix}1&0&\cdots&0&0&\cdots&0\\ 0&1&\cdots&0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{m\times n}$
extracts the first $m$ coordinates of a given $n$-dimensional vector.
Lemma I.1.
For any $x\in{\mathcal{X}}$ and any $\theta\in\mathbb{R}^{n}$,
$\left\langle{x,\theta}\right\rangle=\left\langle{\phi(x),\phi(\theta)}\right\rangle+f(\phi,\theta),$
where $f(\phi,\theta)\in\mathbb{R}$ is some quantity that only depends on $\phi$ and $\theta$ but not $x$.
Proof.
Let $x,x^{\prime}\in{\mathcal{X}}$. By the definition of $\phi$, we have
$\left\langle{\phi(x)-\phi(x^{\prime}),\phi(\theta)}\right\rangle$ $=\left\langle{ZM(x-x^{\prime}),ZM\theta}\right\rangle.$
By the choice of $M$ in Definition 2, $M(x-x^{\prime})$ only has non-zero elements in the first $m$ coordinates. Furthermore, since $Z$ extracts the first $m$ coordinates, we have
$\left\langle{ZM(x-x^{\prime}),ZM\theta}\right\rangle$ $=\sum_{i=1}^{m}(M(x-x^{\prime}))_{i}(M\theta)_{i}$
$=\sum_{i=1}^{n}(M(x-x^{\prime}))_{i}(M\theta)_{i}$
$=\left\langle{M(x-x^{\prime}),M\theta}\right\rangle$
$=\left\langle{x-x^{\prime},\theta}\right\rangle.$ ($M^{\top}=M^{-1}$ because $M$ is a rotation matrix)
Thus,
$\left\langle{x,\theta}\right\rangle-\left\langle{\phi(x),\phi(\theta)}\right\rangle=\left\langle{x^{\prime},\theta}\right\rangle-\left\langle{\phi(x^{\prime}),\phi(\theta)}\right\rangle,$
meaning that the value of $\left\langle{x,\theta}\right\rangle-\left\langle{\phi(x),\phi(\theta)}\right\rangle$ is shared by all $x\in{\mathcal{X}}$. Defining this value as $f(\phi,\theta)$ finishes the proof. ∎
We consider the continuous exponential weight algorithm (Algorithm 5) running on $\phi({\mathcal{X}})\subset\mathbb{R}^{m}$:
1 Let ${\mathcal{X}}\subset\mathbb{R}^{n}$, and let $\phi({\mathcal{X}})\triangleq\{\phi(x):x\in{\mathcal{X}}\}$.
2 for $t=1,2,\ldots$ do
3       Define for $y\in\phi({\mathcal{X}})$,
$w_{t}(y)=\exp\left(\eta\sum_{s=1}^{t-1}\left\langle{y,\phi(\theta_{s})}\right\rangle+\eta\sum_{s=1}^{t}\left\langle{y,\phi(b_{s})}\right\rangle\right)\ \ \text{and}\ \ p_{t}(y)=\frac{w_{t}(y)}{\int_{y^{\prime}\in\phi({\mathcal{X}})}w_{t}(y^{\prime})\mathrm{d}y^{\prime}}$
for some bonus term $b_{t}$.
4      Sample $y_{t}\sim p_{t}$, and play $x_{t}=\phi^{-1}(y_{t})$, where $\phi^{-1}$ is the inverse mapping of $\phi$.
5       Receive $\theta_{t}\in\mathbb{R}^{n}$.
Algorithm 5 Exponential Weights
In Algorithm 5, we require that the inverse mapping of $\phi$ exists. This is true because for any $x,x^{\prime}\in{\mathcal{X}}$, we have $\|\phi(x)-\phi(x^{\prime})\|=\|ZM(x-x^{\prime})\|=\|M(x-x^{\prime})\|=\|x-x^{\prime}\|$, and thus $\phi$ cannot map $x,x^{\prime}\in{\mathcal{X}}$ with $x\neq x^{\prime}$ to the same point.
Theorem I.2.
Let $q_{t}\in\Delta({\mathcal{X}})$ be the distribution such that $x\sim q_{t}$ is equivalent to first drawing $y\sim p_{t}$ and then taking $x=\phi^{-1}(y)$. Algorithm 5 ensures for any $x\in{\mathcal{X}}$,
$\sum_{t=1}^{T}\left\langle{x,\theta_{t}+b_{t}}\right\rangle-\sum_{t=1}^{T}\mathbb{E}_{x\sim q_{t}}\left[\left\langle{x,\theta_{t}+b_{t}}\right\rangle\right]\leq\frac{m\log T}{\eta}+\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{x\sim q_{t}}\left[\exp\left(\eta\left\langle{x,\theta_{t}}\right\rangle\right)-\eta\left\langle{x,\theta_{t}}\right\rangle-1\right].$
Proof.
Note that Algorithm 5 is a standard continuous exponential weight algorithm over reward vectors $\phi(\theta_{t})$ and in the space of $\phi({\mathcal{X}})\subset\mathbb{R}^{m}$. By the standard analysis (see, e.g., Ito et al., (2020); Zimmert and Lattimore, (2022)), we have for any sequence $\lambda_{1},\ldots,\lambda_{T}\in\mathbb{R}$ and any $y\in\phi({\mathcal{X}})$,
$\sum_{t=1}^{T}\left\langle{y,\phi(\theta_{t})+\phi(b_{t})}\right\rangle-\sum_{t=1}^{T}\mathbb{E}_{y\sim p_{t}}\left[\left\langle{y,\phi(\theta_{t})+\phi(b_{t})}\right\rangle\right]$
$\leq\frac{m\log T}{\eta}+\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{y\sim p_{t}}\left[\exp\big{(}\eta\left\langle{y,\phi(\theta_{t})}\right\rangle+\lambda_{t}\big{)}-\big{(}\eta\left\langle{y,\phi(\theta_{t})}\right\rangle+\lambda_{t}\big{)}-1\right].$
By Lemma I.1, the above implies
$\sum_{t=1}^{T}\left\langle{\phi^{-1}(y),\theta_{t}+b_{t}}\right\rangle-\sum_{t=1}^{T}\mathbb{E}_{y\sim p_{t}}\left[\left\langle{\phi^{-1}(y),\theta_{t}+b_{t}}\right\rangle\right]$
$\leq\frac{m\log T}{\eta}+\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{y\sim p_{t}}\left[\exp\Big{(}\eta\left\langle{\phi^{-1}(y),\theta_{t}}\right\rangle-\eta f(\phi,\theta_{t})+\lambda_{t}\Big{)}-\Big{(}\eta\left\langle{\phi^{-1}(y),\theta_{t}}\right\rangle-\eta f(\phi,\theta_{t})+\lambda_{t}\Big{)}-1\right],$
which further implies that for any $x\in{\mathcal{X}}$,
$\sum_{t=1}^{T}\left\langle{x,\theta_{t}+b_{t}}\right\rangle-\sum_{t=1}^{T}\mathbb{E}_{x\sim q_{t}}\left[\left\langle{x,\theta_{t}+b_{t}}\right\rangle\right]\leq\frac{m\log T}{\eta}+\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{x\sim q_{t}}\left[\exp\left(\eta\left\langle{x,\theta_{t}}\right\rangle\right)-\eta\left\langle{x,\theta_{t}}\right\rangle-1\right]$
by the definition of $q_{t}$ and by letting $\lambda_{t}=\eta f(\phi,\theta_{t})$. ∎
Appendix J Computationally Efficient Algorithm for Adversarial $C$ Bound